OpenStudy (anonymous):

Find the slope-intercept form of the equation of a line with slope -2 containing point (2, -9). Question 7 options: y = 2x - 5 y = -2x - 13 y = -2x - 5 y = -5x + -2

hero (hero):

y - y1 =m(x - x1) y + 9 =-2(x - 2) y = -2x + 4 - 9 y = -2x - 5

OpenStudy (anonymous):

thanks